If a random sample of size 81 is drawn from this

population, what is the probability that the mean

of this sample is larger than 53.5 inches?


1
Expert's answer
2022-05-03T16:11:05-0400

Let "X=" the sample mean: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=54, \sigma=2.5, n=81."


"P(X>53.5)=1-P(Z\le\dfrac{53.5-54}{2.5/\sqrt{81}})"

"=1-P(Z\le-1.8)\approx0.96407"

"96.407\%"